OpenStudy (anonymous):

An isosceles triangle where the ratio of the vertex angle to one of the base angles is 1:2. Find the measures of each angle in the triangle.

OpenStudy (welshfella):

the ratio of 3 angles = 1:2:2 so the vertex angle = 1/(2+2+1) * 180 = 1/5 * 180 and the base angles are both twice that result